# Fast fitting of phylogenetic mixed effects models

### v1

@@ Introduction
With this choice of a matrix normal covariance strucutre, VA can still be slow due to the potentially large number of variational parameters in $\textbf{A}_{m}$. The number of variational parameters in this matrix grows quadratically in the number of species. To ensure better scaling, and quick computation also with many species, we employ a third approximation by representing the off-diagonal entries of $\textbf{A}_{m}$ in a reduced-rank fashion so that $d<<m$, i.e., $\textbf{A}_{m}=\textbf{A}_{d}\textbf{A}_{d}^{\top}+\textbf{A}_{m-d}$ where $\textbf{A}_{d}$ is a $m\times d$ lower triangular matrix with positive diagonal entries, and $\textbf{A}_{m-d}$ is a $m\times m$ diagonal matrix with the first $d$ diagonal entries equal to zero and the remaining entries positive, ensuring that $\textbf{A}_{m}$ is of full rank. Such a choice for reducing the number of variational parameters in the likelihood is optimal for computation speed, to reduce memory use during model compilation, and to reduce the size of the Hessian matrix that needs to be inverted for parameter estimation and for post-hoc estimating (approximate) statistical uncertainties. This does introduce the additional problem of having to select the rank of the variational covariance matrix, but in the simulation studies that follow we show that high accuracy can be achieved even when using only a single dimension.
@@
The final remaining computational bottleneck involves inversion of the phylogenetic covariance matrix. Since the covariance matrix for the covariate effects $\boldsymbol{\Sigma}_{r}$ is usually relatively small, its inverse can be evaluated explicitly without excessive computational cost, and its determinant retrieved by formulating it following a log-cholesky parameterization (Pinheiro and Bates 1996). In contrast, the number of species in multivariate ecological data is often relatively large, so that $\boldsymbol{\Sigma}^{-1}$ can be expensive to calculate. With a single phylogenetic signal parameter over all covariate effects, the form of the covariance function for $\boldsymbol{\Sigma}$ allows inversion by repeatedly applying the Sherman-Morrison formula for computationally cheap rank-one updates to the inverse of C, but this still has $O(m^{2})$ computational complexity, so further improvement is warranted. Recently, (Matsuba et al. 2024) employed a nearest neighbor Gaussian Process (NNGP) approximation to quickly and accurately retrieve the inverse of the phylogenetic matrix, although with a different model parameterisation than that we use here. We also adopt NNGPs as our fourth and final approximation here.
We do not discuss the technical details of NNGPs in-depth here, but instead refer to Finley et al. (2019) for further reading on the matter. In brief, the joint distribution of the random effects per covariate can be written as the product of conditional distributions $p(\boldsymbol{\beta}_{k})=\prod^{m}_{j=1}p(\beta_{1k})p(\beta_{2k}|\beta_{1k}),\ldots,p(\beta_{mk}|\beta_{1k},\ldots,\beta_{j-1k})$. Rather than conditioning on the effects of all previous species, in NNGPs we instead reduce the size of the set of species we condition on by using the set of $l$ nearest neighbours. Due to the conditional independence assumption that is introduced, this results in a sparse approximation of the phylogenetic precision matrix that is usually very accurate even for a small number of nearest neighbors. NNGPs facilitate parallel computation well, unlike regular matrix inversion that tends to be inherently sequential (such as repeated application of the Sherman-Morrison formula), because the calculation can be performed for all species separately, allowing us to somewhat leverage the benefits of sparse matrix algebra routines, and to better parallize model fitting. Instead of directly constructing the sparse precision matrix, in NNGPs we instead construct the (approximate) inverse of the cholesky factor of the covariance matrix. This combines nicely with the likelihood in equation (5), since that only depends on the phylogenetic precision matrix, and it makes for an efficient implementation of equation (3), as there is no extra cost for calculating the cholesky factorization of the phylogenetic precision matrix per covariate effect.
@@
For the first simulation study, we kept the number of sites fixed at $n=100$ and the number of species at $m=200$. We simulated 250 datasets while varying the rank of the VA covariance matrix ($q\in\{1,3,5,10,25,100,200\}$). Instead of the NNGP, in this first simulation study we calculated the inverse of the phylogenetic covariance matrix by repeated rank-1 updates, for as quick as possible inversion, but without introducing potential error due to the NNGP so that we may isolate the error solely due to the low-rank VA approximation. For comparison we additionally fitted the model with the Laplace approximation as automatically applied by Template Model Builder (Kristensen et al. 2015). We re-simulated the phylogeny for each simulation using the ape R-package (Paradis and Schliep 2019). We simulated five covariates from $\text{Uniform}(-1,1)$, the covariance matrix for the random effects from $W_{p}(\frac{1}{5}\textbf{I}_{p},p)$, and kept $\rho=0.5$ fixed so that the median absolute error is at best zero and at most 0.5. As the total number of simulations is high, and the models take long to fit when the rank of the VA approximation is large, we only simulated binary datasets. Binary datasets contain least information of the datatypes that are accommodated by the gllvm R-package, so that these simulations should provide a reliable lower bound to the accuracy of the approximation. The results of the simulations are summarized in Figure 1.
@@
Figure 1: Results of the first simulation study for evaluating the accuracy of the low-rank VA approximation. The left panel displays increase of fitting time with the rank (1, 3, 5, 10, 25, 100, 200) of the VA approximation, where the dashed red line indicates the computation time for the models fitted by Laplace approximation, and the dashed grey lines the 2.5% and 97.5% percentiles. The right panel shows the median absolute error for the phylogenetic signal parameter, which at most can be 0.5. Again, the red line indicates performance of the Laplace approximation and the grey dashed lines the 2.5% and 97.5% percentiles.
@@
The first simulation study confirms that the computation time of the models rapidly increase with the selected rank of the VA approximation, even far beyond the fitting time for the Laplace approximation (indicated by the red dashed line). As seen in the left panel of Figure 1, the rank of the VA approximation had a large impact of the computation time, but fortunately little impact on the estimation accuracy of the phylogenetic signal parameter, which remained nearly stable across various ranks of the VA approximation and was indeed lowest for the full rank approximation.
@@
In the second simulation study we kept the rank of the VA approximation fixed to one, as well as the number of sites fixed to $n=100$, but instead varied the number of species ($m\in\{50,100,250,500\}$) and the number of nearest neighbors used ($nn\in\{3,5,7,9,11,13,15\}$) to approximate the inverse of the phylogenetic covariance matrix. For each combination we simulated 250 datasets, for either binary data or counts. The results are presented in Figure 2.
@@
The results in Figure 2 represent the compounded approximation error due to both the low-rank VA approximation and the NNGP. The number of species has a larger impact on the estimation accuracy of the phylogenetic signal than the NNGP approximation. As the number of species increase, the estimation accuracy of the phylogenetic signal improved, with $m=50$ being too few species for accurate estimation, but with $m>100$ usually providing accurate estimates. The accuracy was higher with more nearest neighbours, but a relatively low (e.g., 10) number of nearest neighbors seems sufficient for accurate estimates of the phylogenetic signal, even for large numbers of species in the data.
@@
Here fit the same four models: three fourth corner models with 1) fruit-body type, (log) spore volume, and trophic group (“lifestyle”; wood decaying or other), 2) Principal Component axes extracted from the matrix of all traits, 3) the first axis due to each a) the reproductive traits, b) dispersal traits, and c) resource use traits, and 4) a model without traits. All models included covariate-specific phylogenetic signal parameters, except for one additional model without traits that was fitted for better comparison with the results from Abrego et al. (2022a). We decided against incorporating latent variables for brevity, and because the gllvm R-package at the time of writing only facilitates including latent variables at the observation level.
In the Hmsc R-package each model takes about a week to 10 days to run (N. Abrego, pers. comm., 1st March 2024), while it takes three to four minutes per model with the methods in this article (on a Dell Latitude 7490 using 7 CPU). This was with 10 nearest neighbors, nested random intercepts for the study region and reserves, and including correlation between covariate effects.
The phylogenetic signal parameter estimates from all models are organised in table 1. Our results deviated somewhat from the original analysis, in that the phylogenetic signal was estimated to be somewhat lower, with 0.51 being the estimated phylogenetic signal for a model without traits, and while assuming the phylogenetic signal to be the same for all covariates, compared to the 0.73 found in the original analysis by Abrego et al. (2022a).
@@
Table 1: Phylogenetic signal parameter estimates for models fitted to the Abrego et al. (2022) data. A parameter estimate near zero is displayed as a blank cell for improved readability.
No traits(a) No traits(b) Group-specific PCA Raw traits Overall PCA
Intercept 0.51 0.04 0.04 0.02 0.05
Deadwood size 0.51 0.12 0.01 0.01 0.01
Decay stage 0.51 0.99 0.94 0.85 0.91
Decay stage² 0.51 0.82 0.32 0.18 0.10
Connectivity 0.51 0.73 0.59 0.49 0.47
Temperature range 0.51 0.06 0.02
Precipitation 0.51 0.69 0.64 0.57 0.68
ln(Reserve area) 0.51 0.43 0.39 0.42 0.53
The Hmsc R-package is presently unable to estimate covariate-specific phylogenetic signal parameters, so that Abrego et al. (2022a) post-hoc retrieved estimates of phylogenetic signal by fitting phylogenetic regressions to the posterior distributions of the covariate effects, using the ape R-package. At the deadwood-unit level, they found species responses to decay stage, and decay stage² to be phylogenetically structured, and at the forest-site level they found reserve area, and in some models connectivity, to be phylogenetically structured. Abrego et al. (2022a) did not find a statistically significant result for the effect of precipitation.
At the deadwood-unit level, we similarly find species responses to decay stage to be phylogenetically structured in all of the fitted models, and when traits are absent also to decay stage². At the forest-site level, we are able to corroborate the result by Abrego et al. (2022a) that species responses to connectivity and reserve area are phylogenetically structured to some degree, though we additionally find species responses to precipitation to be phylogenetically structured. Consequently, although the exact numerical values differ, we find similar phylogenetic structuring at both spatial scales present in the study, but while fitting the model in a single step and much more quickly.
For further visual presentation of the results we have chosen the model with the original trait variables. This was ranked lowest by AIC, and provides the most straightforward interpretation (unlike the models that include Principal Component axes). This result confirms that species responses were indeed structured by functional traits. Figure 3 displays the phylogeney with the predicted phylogenetic random effects, as well as the community mean responses, and Figure 4 displays the joint effects of the community mean response and the effect of trait on species responses, with their 95% confidence interval bands (see appendix S3 for a visualization of the marginal trait-environment effects).
@@
Figure 3: The left panel displays community mean responses to the covariates, with 95% confidence intervals. Coefficients are greyed out of the confidence interval crosses zero, so that it is unclear if there is an effect of a covariate on the community. The bottom panel presents species-specific deviations from the community mean effects. A species-specific effect of zero is interpreted as lack of evidence that a species responds differently to the environment than the community mean response. Orange colors indicate a more negative response than the community mean effect, and green more positive. Species effects of which the prediction interval included zero have been removed.
For the intercept, dead wood size, and decay stage² there was sufficient evidence to conclude a non-zero community mean response. The estimate for decay stage² was negative, and predicted species-specific random effects generally smaller than the community mean response, corresponding to typically expected unimodal (bell-shaped) curves that open downward (i.e., with maxima).
Species responses to deadwood size were positive on average, but become increasly less positive with an increase in spore volume, with Corticioid fruit-body types having the least positive response and Polyporoid pileate fruit-body types the most positive. The effect of decay stage is difficult to disentangle due to the additional quadratic term in the model, though the effect of the quadratic term can be used to infer how species tolerances change with traits (van der Veen et al. 2021). Polyporoid resupinate and Polyporoid pileate fruit-body types exhibit a more narrow tolerance to changes in decay stage than Stromatoid fruit-body types, i.e., they are more sensitive to change in decay stage of deadwood.
Wood decaying fungi were most positively affected by connectivity of the forest among fungi species, and by the size of the reserve. Fungi generally exhibited increased presence to improved connectivity of the forest, and wood-decaying fungi with large spore volume especially so. Fungi with larger spore volume exhibited a more negative response to precipitation than species with a small spore volume. Finally, wood-decaying fungi are among the most prevalent fungi types in larger reserves.
Figure 4: Predicted species responses to the environment from the model as a function of traits, with 95% confidence interval bands. Grey points represent the estimated species responses to the environmental covariates. Different fruit-body types are represented as different color lines, whether fungi are wood-decaying as dashed lines, and the final trait of spore volume is visualized on the horizontal axes

### v2

@@ Introduction
With this choice of a matrix normal covariance structure, VA can still be slow due to the potentially large number of variational parameters in $\textbf{A}^{m}$. The number of variational parameters in this matrix grows quadratically in the number of species. To ensure better scaling, and quick computation also with many species, we employ a third approximation by representing the off-diagonal entries of $\textbf{A}^{m}$ in a reduced-rank fashion so that $d<<m$, i.e., $\textbf{A}^{m}=\textbf{A}^{d}\textbf{A}^{d\top}+\textbf{A}^{m-d}$ where $\textbf{A}^{d}$ is a $m\times d$ lower triangular matrix with positive diagonal entries, and $\textbf{A}^{m-d}$ is a $m\times m$ diagonal matrix with the first $d$ diagonal entries equal to zero and the remaining entries positive, ensuring that $\textbf{A}^{m}$ is of full rank. Such a choice for reducing the number of variational parameters in the likelihood is optimal for computation speed, to reduce memory use during model compilation, and to reduce the size of the Hessian matrix that needs to be inverted for parameter estimation and for post-hoc estimating (approximate) statistical uncertainties. This does introduce the additional problem of having to select the rank of the variational covariance matrix, but in the simulation studies that follow we show that high accuracy can be achieved even when using only a single dimension.
Sparse precision matrix
@@
The final remaining computational bottleneck involves inversion of the covariance matrix for the random effects. Since the covariance matrix of the covariate effects $\boldsymbol{\Sigma}^{r}$ is usually relatively small, its inverse can be evaluated explicitly without excessive computational cost, and its determinant straightforwardly retrieved by formulating it following a log-cholesky parameterization (Pinheiro and Bates 1996). In contrast, the number of species in multivariate ecological data is often relatively large, so that $\boldsymbol{\Sigma}^{-1}_{k}$ is expensive to calculate. With a single phylogenetic signal parameter over all covariate effects, the form of the covariance function for $\boldsymbol{\Sigma}_{k}$ allows inversion by repeatedly applying the Sherman-Morrison formula for computationally cheap rank-one updates of the form $\textbf{C}^{-1}=\textbf{C}^{-1}-\{1-\rho\}/\{1+(1-\rho)\textbf{C}_{j}\textbf{C}_{j}^{\top}\}$, but further improvement is warranted.
Let $A_{j}$ a set of indices (species) used to condition a random effect on. Consequently, the joint distribution for the random effect over all species can be written as $p(\boldsymbol{\beta}^{\epsilon}_{k})=\prod\limits^{j=1}p(\beta^{\epsilon}_{jk}|\{\beta^{\epsilon}_{1k},\beta^{\epsilon}_{2k},\ldots,\beta^{\epsilon}_{j-1k}\}:A_{j})$, and while momentarily ignoring potential correlations between random effects, each conditional distribution can be written as $p(\beta^{\epsilon}_{jk}|\{\beta^{\epsilon}_{1k},\beta^{\epsilon}_{2l},\ldots,\beta^{\epsilon}_{j-1k}\}:A_{j})=\mathcal{N}(\boldsymbol{\Sigma}_{kjA_{j}}\boldsymbol{\Sigma}_{kA_{j}A_{j}}^{-1}\boldsymbol{\beta}^{\epsilon}_{A_{j}k},\boldsymbol{\Sigma}_{kjj}-\boldsymbol{\Sigma}_{kjA_{j}}\boldsymbol{\Sigma}_{kA_{j}A_{j}}^{-1}\boldsymbol{\Sigma}_{kA_{j}j})$.
Wist and Rue (2006) describe various methods for specifying Gaussian Markov Random Fields with a sparse precision matrix, such as by thresholding or by a predefined sparsity pattern (such as a band pattern). Here, we mostly focus on adopting NNGPs as our fourth and final approximation, but compare with approximating the precision matrix by a band pattern in the real data example. Unlike regular matrix inversion, that tends to be inherently sequential (such as repeated application of the Sherman-Morrison formula), NNGPs facilitate parallel computation well because the calculation can be performed for all random effects separately. In NNGPs we construct the (approximate) inverse of the cholesky factor of the covariance matrix, which makes for an efficient implementation of equation (2), and combines nicely with the likelihood in equation (3).
NNGPs are usually applied in a spatial (and Bayesian) context (Finley et al. 2019), truncating the set of indices $A_{j}$ for spatial locations based on geographical proximity. For a phylogenetic random effect we instead condition each species’ random effect on the species closest in evolutionary space. Distance between species in evolutionary space is readily accessible from the phylogeny, and calculated as the total shared branch length on the phylogeny. It is this truncation for the conditional distributions of the random effects that implies a sparse precision matrix for the joint distribution of the random effects. Note, NNGPs are order-sensitive, so that the approximation can suffer from a poor ordering of species. However, similarly to Guinness (2018), we choose to see this as a feature of the methodology, so that the approximation can be improved by permuting the columns of the response data. We further explore this issue of order sensitivity in the real data example.
@@
For the first simulation study, we kept the number of sites fixed at $n=100$ and the number of species at $m=200$. We simulated 250 datasets while varying the rank of the VA covariance matrix ($d\in\{1,3,5,10,25,100,200\}$). Instead of the NNGP, in this first simulation study we directly calculated the inverse of the phylogenetic covariance, in order to avoid introducing potential error due to the NNGP and so that we may isolate the error solely due to the matrix normal VA approximation with reduced rank, and unstructured or diagonal $\textbf{A}^{r}$. For comparison we additionally fitted the model with the Laplace approximation as automatically applied by Template Model Builder (TMB, Kristensen et al. 2015). We re-simulated the phylogeny for each simulation using the ape R-package (Paradis and Schliep 2019). We simulated five covariates from $\text{Uniform}(-1,1)$, the covariance matrix for the random effects from $W(\frac{1}{5}\textbf{I}^{p},p)$, and kept $\rho=0.5$ fixed so that the median absolute error is at best zero and at most 0.5. As the total number of simulations is high, and the models take long to fit when the rank of the VA approximation is large, we only simulated binary datasets. Binary datasets contain least information of the datatypes that are accommodated by the gllvm R-package, so that these simulations should provide a reliable upper bound to the accuracy of the approximation. The results of the simulations are summarized in Figure 1.
@@
Figure 1: Results of the first simulation study for evaluating the accuracy of the low-rank VA approximation. The left panel displays increase of fitting time with the rank (1, 3, 5, 10, 25, 100, 200) of the VA approximation, with solid black lines indicating models fitted with the VA approximation and with unstructured $\textbf{A}^{r}$ (circles) or diagonal $\textbf{A}^{r}$ (triangles), while the dashed red line indicates the computation time for the models fitted by Laplace approximation. The faded black lines represent the 2.5% and 97.5% percentiles for the VA approximation and the faded dashed red lines for the Laplace approximation. The right panel shows the median absolute error for the phylogenetic signal parameter, which at most can be 0.5.
@@
The first simulation study confirms that the computation time of the models rapidly increase with the selected rank of $\textbf{A}^{m}$ in the VA approximation, while being competitive to the Laplace approximation implemented with TMB (indicated by the red dashed line). As seen in the left panel of Figure 1, the rank of the VA approximation has a large impact of the computation time, but fortunately little impact on the estimation accuracy of the phylogenetic signal parameter, which remained nearly stable across various ranks of the VA approximation and was indeed lowest for the full rank approximation. Models fitted with an unstructured $\textbf{A}^{r}$ in the VA approximation exhibited longer fitting times, so that it took longer to fit models than with the Laplace approximation. Models fitted with a diagonal $\textbf{A}^{r}$ fitted more quickly than with the Laplace approximation, while also providing more accurate estimates for the phylogenetic signal than with an unstructured $\textbf{A}^{r}$.
@@
In the second simulation study we kept the rank of the VA approximation fixed to one and $\textbf{A}^{r}$ unstructured, and fixed the number of sites to $n=100$, but instead varied the number of species ($m\in\{50,100,250,500\}$) and the number of nearest neighbors used ($nn\in\{1,3,5,7,9,11,13,15\}$) to approximate the inverse of the phylogenetic covariance matrix. For each combination we simulated 250 datasets, for either binary data or counts. The results are presented in Figure 2.
@@
The results in Figure 2 represent the compounded approximation error due to both the low-rank VA approximation and the NNGP. The number of species has a larger impact on the estimation accuracy of the phylogenetic signal than the NNGP approximation. As the number of species increases, the estimation accuracy of the phylogenetic signal improves, with $m=50$ potentially requiring a much larger number of nearest neighbouurs for accurate estimation, but with $m>100$ providing accurate estimates under most scenarios. As expected, the accuracy was higher with more nearest neighbours, but a relatively low (e.g., 10) number of nearest neighbors seems sufficient for reasonably accurate estimates of the phylogenetic signal, even for large numbers of species in the data. Though not presented here, application of NNGPs reduced computation time significantly relative to the models fitted for the first simulation study.
@@
The models fitted by Abrego et al. (2022a) additionally incorporated six latent variables at the region-level, and although the gllvm R-package facilitates including such latent variables at the group-level, we have here chosen to omit the latent variables for brevity. In the Hmsc R-package each model took about a week to 10 days to run (N. Abrego, pers. comm., 1st March 2024). Results for the same four models as in are included in appendix S3. It takes about six to twenty minutes to fit a model with covariate-specific phylogenetic signal parameters (on a Dell Latitude 7490 using 7 CPU, with $nn=15$ neighbours, and with unstructured $\textbf{A}^{r}$), and two to four minutes for models while assuming $\rho_{k}=\rho$, i.e., a single phylogenetic parameter for all covariate effects. Fitting the models with diagonal $\textbf{A}^{r}$ can further reduce computation time, to (e.g.) 1.5 minute (corresponding to the two minutes for a model without traits). These models incorporated nested random intercepts for the study region and reserves, and including all correlation parameters between the covariate effects.
As previously noted, the quality of the NNGP approximation tends to be sensitive to the order of species. This has previously been noted by Guinness (2018) in a spatial context, who also explored orderings for improving the NNGP approximation. Similarly to finding permutation methods for reducing fill-in of the cholesky factor of a sparse matrix (Wist and Rue 2006), identifying an optimal ordering for NNGPs is an NP-complete problem. Consequently, heuristic methods are essential for determining an ordering that improves the NNGP approximation.
The approximation error for the phylogenetic precision matrix will have to be assessed on a case-by-case basis, but can be explored before model fitting, because the approximation error for $\boldsymbol{\Sigma}_{k}$ is clearly related to that of C. However, the former is not available before model fitting, whereas the approximation to $\textbf{C}^{-1}$ is. We define the approximation error as $\|\widetilde{\textbf{C}^{-1}}\textbf{C}-\textbf{I}^{m}\|_{F}$, where $\widetilde{\textbf{C}^{-1}}$ denotes the approximation to $\textbf{C}^{-1}$.
Rather than extensively searching for potentially suitable orderings, we instead focus on exploring the approximation error due to six orderings that are easily accessible to ecologists: 1) the order of species in the phylogeny; because species with large covariances are naturally located closely together in the phylogeny, 2) alphabetically; as a baseline and because it is how data is often provided in practice, 3) the sum of pairwise distances of species, 4) the distance of species to the root of the phylogeny, 5) by the first eigenvector of C (as in Matsuba et al. 2024), 6) by the sum of squared covariances. We additionally explore the approximation error using a band pattern (as in Wist and Rue 2006), ignoring the distance between species in the phylogeny. The results are visualized in Figure 3.
Figure 3: Approximation error for the band pattern approximation and the NNGP, for different permutations of species in the fungi data.
The results show that the NNGP has the potential of a smaller approximation error, though it requires conditioning on six or more species. When conditioning on less than six species, the pre-defined band pattern approximation exhibits a smaller approximation error. For the band pattern approximation, ordering the response data following the order of species in the phylogeny provides the smallest approximation error, while for the NNGP this is when ordering the response data by the total shared branch length in the phylogeny. In both instances, ordering species alphabetically made for the worst approximation, even though as a result species within the same genus will usually be located in close proximity (which can be expected to be phylogenetically similar, and thus have large covariances). The same does not hold for genera that are phylogenetically similar.
Figure 4: The phylogenetic covariance matrix as a result from the sparse approximation to the precision matrix for different orderings and numbers of nearest neighbours (nn). Three orderings are displayed, corresponding to poor (alphabetical), average (tips) and good (distance) approximations.
The approximation error affects both the estimation accuracy and computation speed. To demonstrate this, we fitted three models to the example data using three different orderings and $nn\in\{5,10,15,20\}$, the results are displayed in Table 1. The models included three trait covariates, all environmental covariates, random intercepts for region and reserve. We also assumed a diagonal $\textbf{A}^{r}$ for faster fitting. In Figure 5 the community mean responses and species’ random effects for the model with $nn=20$ with distance ordering are displayed alongside the phylogeny.
@@
Table 1: Phylogenetic signal parameter estimates for models with different orderings of species in the response data, and number of nearest neighbours (nn), fitted to the Abrego et al. (2022) data. A parameter estimate of 0.05 or lower is displayed as a blank cell for improved readability.
Order nn Time (min) Intercept Deadwood size Decay stage Decay stage² Connectivity Temperature Precipitation ln(Reserve area)
alphabetical 5 8.77 0.13 1.00 0.88 0.13 0.51 0.55
alphabetical 10 5.92 0.24 0.65 1.00 0.87 0.10 0.74 0.68
alphabetical 15 5.96 0.22 0.66 1.00 0.88 0.14 0.71 0.70
alphabetical 20 9.33 0.22 0.64 1.00 0.88 0.15 0.64 0.71
tips 5 11.98 0.17 0.62 0.99 0.89 0.10
tips 10 6.96 0.07 0.98 0.89 0.52 0.74
tips 15 7.81 0.49 0.99 0.90 0.59 0.77
tips 20 8.07 0.50 0.99 0.91 0.59 0.81 0.11
distance 5 5.63 0.18 0.68 0.88 0.12 0.10 0.16
distance 10 6.02 0.26 0.99 0.86 0.07
distance 15 8.74 0.31 0.99 0.85
distance 20 6.07 0.31 0.98 0.87 0.19 0.46
@@
Figure 5: The right panel displays community mean responses to the covariates, with 95% confidence intervals. Coefficients are greyed out if the confidence interval crosses zero, so that it is unclear if there is an effect of a covariate on the community. The bottom panel presents species-specific deviations from the community mean effects. A species-specific effect of zero is interpreted as lack of evidence that a species responds differently to the environment than the community mean response. Orange colors indicate a more negative response than the community mean effect, and green more positive. Species effects of which the prediction interval included zero have been removed.